Linear Equation from slope and point: Suppose that f is a linear function with slope 2.6 and that f(5)=-3. What value of x gives f(x)=0?
2nd Question: Linear Equation from two points: Suppose f is a linear function such that f(3)=5 anf f(7)=-4. Find the equation for f.

Thank you, Andie

Answer by stanbon(75887)

Linear Equation from slope and point: Suppose that f is a linear function with slope 2.6 and that f(5)=-3. ------
y = mx + b
-3 = 2.6*5 + b
-16 = b
Equation:
f(x) = 2.6x -16
-------
What value of x gives f(x)=0?
0 = 2.6x - 16
x = 6.1538....
==================

2nd Question: Linear Equation from two points: Suppose f is a linear function such that f(3)=5 anf f(7)=-4. Find the equation for f.
-----
slope = (-4-5) / (7-3) = -9/4
intercept: 5 = (-9/4)*3 + b
b = 5+(27/4) = 47/4
Equation:'
y = (-9/4)x + (47/4)
==========================
Cheers,
Stan H.